OpenStudy (anonymous):

2/3 (x - 2) = 4x. slove for x

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

So basically i would do 2/3 X x/-2 ?

OpenStudy (anonymous):

1-(1/6)√6√7,(1/6)√6√7+1

OpenStudy (anonymous):

?

OpenStudy (phi):

is the (x-2) in the denominator or is it \[ \frac{2}{3} (x-2) \]

OpenStudy (anonymous):

Um i'm not sure

OpenStudy (phi):

is it \[ \frac{2}{3(x-2)} = 4x \] or \[ \frac{2}{3}(x-2)= 4x \]